# Self-Paced and Self-Corrective Masked Prediction for Movie Trailer Generation

### v1

@@ 4.2 Comparisons
To evaluate the perceptual quality of the generated trailers, we conduct a user study involving 25 participants. Each participant compares the trailers generated by our method and several baselines with the corresponding official trailers, where all trailers are anonymous to ensure unbiased evaluation. Following in the work in [19, 50, 54], the evaluation is performed from four aspects: theme (i.e., how well the trailer conveys the theme of the movie), rhythm (i.e., how well the visuals align with the rhythm of the music), attractiveness (i.e., how attractive the trailer appears to the viewer), and appropriateness (i.e., how close the trailer is to the official trailer), which together provide a comprehensive assessment of the perceived gap between generated and official trailers. The score of each aspect ranges from 1 to 7, with higher scores corresponding to better performance. The radar chart in Fig. 3(a) illustrates the average scores of all participants and demonstrates that our method consistently outperforms previous methods.
@@ 4.3 Ablation Studies
Shot Deviation. Referring to the work in [1, 50, 44], we also conduct experiments on allowable positional deviation during trailer generation. This design is motivated by the observation that adjacent shots in movies often belong to the same event and exhibit highly similar visual content, making them difficult to distinguish precisely. To mitigate the potential bias caused by such local ambiguity and ensure a more reasonable evaluation of model performance, we relax the matching criterion by including the $R$ neighboring shots around each official trailer shot within the ground truth set. The results in Tab. 7 show that allowing small positional deviations consistently improves performance, confirming that adjacent shots typically share overlapping semantics and that strict positional alignment is therefore not essential for assessing model performance.
@@ 6 Model Architecture
where $R_{\theta}$ represents the rotary positional embedding matrix [42] to encode relative positions, and $W_{Q}$, $W_{K}$ and $W_{V}$ are learnable matrices. The feed-forward network (FFN) uses a gated SiLU activation with a linear expansion and projection:
@@
where $\text{Linear}_{\text{up}}$: $\mathbb{R}^{D}\rightarrow\mathbb{R}^{2D}$, and $\text{Linear}_{\text{out}}$: $\mathbb{R}^{2D}\rightarrow\mathbb{R}^{D}$. Referring to [33], our model replaces standard layer normalization with RMS normalization:

### v2

@@ 4.2 Comparisons
To evaluate the perceptual quality of the generated trailers, we conduct a user study involving 25 participants. Each participant compares the trailers generated by our method and several baselines with the corresponding official trailers, where all trailers are anonymous to ensure unbiased evaluation. Following in the work in [irie2010automatic, wang2024inverse, zhu2025weakly], the evaluation is performed from four aspects: theme (i.e., how well the trailer conveys the theme of the movie), rhythm (i.e., how well the visuals align with the rhythm of the music), attractiveness (i.e., how attractive the trailer appears to the viewer), and appropriateness (i.e., how close the trailer is to the official trailer), which together provide a comprehensive assessment of the perceived gap between generated and official trailers. The score of each aspect ranges from 1 to 7, with higher scores corresponding to better performance. The radar chart in Fig. 3(a) illustrates the average scores of all participants and demonstrates that our method consistently outperforms previous methods.
@@ 4.3 Ablation Studies
Shot Deviation. Referring to the work in [argaw2024towards, wang2024inverse, sun2025multi], we also conduct experiments on allowable positional deviation during trailer generation. This design is motivated by the observation that adjacent shots in movies often belong to the same event and exhibit highly similar visual content, making them difficult to distinguish precisely. To mitigate the potential bias caused by such local ambiguity and ensure a more reasonable evaluation of model performance, we relax the matching criterion by including the $R$ neighboring shots around each official trailer shot within the ground truth set. The results in Tab. 7 show that allowing small positional deviations consistently improves performance, confirming that adjacent shots typically share overlapping semantics and that strict positional alignment is therefore not essential for assessing model performance.
@@ 6 Model Architecture
where $R_{\theta}$ represents the rotary positional embedding matrix [su2024roformer] to encode relative positions, and $W_{Q}$, $W_{K}$ and $W_{V}$ are learnable matrices. The feed-forward network (FFN) uses a gated SiLU activation with a linear expansion and projection:
@@
where $\text{Linear}_{\text{up}}$: $\mathbb{R}^{D}\rightarrow\mathbb{R}^{2D}$, and $\text{Linear}_{\text{out}}$: $\mathbb{R}^{2D}\rightarrow\mathbb{R}^{D}$. Referring to [nie2025large], our model replaces standard layer normalization with RMS normalization: